OpenStudy (anonymous):

12 people is what percent of 48 people?

OpenStudy (anonymous):

ill give medal

OpenStudy (redheadangel):

25%

OpenStudy (anonymous):

so how did you get that answer

OpenStudy (redheadangel):

12 divided by 48 :)

OpenStudy (redheadangel):

that gets you .25 which is equivalent to 25%

OpenStudy (anonymous):

percent problems can be easily solved by using the percent formula : is/of = %/100 Now look at your problem and assign values " is " = 12 " of " = 48 % = x (because this is what we are solving for now we sub into percent formula 12/48 = x/100 --- since this is a proportion, we must cross multiply (48)(x) = (12)(100) 48x = 1200 -- divide both sides by 48 x = 25 so...12 people is 25% of 48 you can use this formula for any percent problem...just be sure and assign the numbers to the right position

OpenStudy (anonymous):

thanks